OpenStudy (anonymous):

Which ordered pair is the solution to the system of equations? A. (−3, −2) B. (−1, 3) C. (1, 5) D. (−4, 0)

OpenStudy (anonymous):

can someone plz help ASAP i really need it, i only have 5 questions i will medal whoever comes to help me first but i have to see what grade i get to medal u.

OpenStudy (anonymous):

hello anyone

OpenStudy (ramirocruzo):

Hmm....Where are the system of equations @savana24 ??

OpenStudy (anonymous):

sorry hold on

OpenStudy (anonymous):

it doesn't have it i copied and pasted everything that was on the question except what number i am on

OpenStudy (anonymous):

plz hurry i have to go to bed right now and i still have a couple more tests

OpenStudy (ramirocruzo):

We cant help until you give us the equations...Send a screenshot..

OpenStudy (anonymous):

didn't u hear me there is no equations to the question that is the only thing that they gave me

OpenStudy (anonymous):

i think these are it 2x+y = -8 x-y = -4

OpenStudy (anonymous):

yea they are i am so sorry i had copied and pasted it but those equations didn't paste on what i gave u idk why so sorry.

OpenStudy (anonymous):

and i saw that the equations were blue

OpenStudy (ramirocruzo):

Its D...Multiply second eqn. with two then subtract its from first eqn...

OpenStudy (anonymous):

Which ordered pair is the solution to the system of equations? A. (2, −1) B. (1, −9) C. (−2, −2) D. (−6, −3)

OpenStudy (anonymous):

the equations to this one is: X-4Y =6 3X+Y = -8

OpenStudy (anonymous):

plz hurry my mom is coming to ask what i have done and this is the only one i have done just give me the answers i will fan u.

OpenStudy (ramirocruzo):

Hmm....Its C

OpenStudy (anonymous):

Which ordered pair is the solution to the system of equations? A. (8, 1) B. (−12, 2) C. (−4, −8) D. (4, −2)

OpenStudy (anonymous):

the equations for this one is: -3X + 4Y = -20 X+4Y = -4

OpenStudy (nincompoop):

you can evaluate the systems of equation by either substitution or elimination method

OpenStudy (triciaal):